# RDD: Robust Feature Detector and Descriptor using Deformable Transformer

### v4

@@ 3.3 Feature Matching
Dual-Softmax operator [35, 44] is used to establish correspondences between two descriptors $d^{1,2}$. The score matrix $S$ between the descriptors is first calculated by $S\left(i,j\right)=\frac{1}{\tau}\cdot\langle\ d^{1},d^{2}\rangle$, where $\tau$ is the temperature. We can now apply softmax on both dimensions of $S$ to obtain the probability of soft mutual nearest neighbor (MNN) matching. The matching probability matrix $P$ is obtained by:
@@
Recent works [42, 5, 17, 46] demonstrated the benefits of semi-dense feature matching by improving coverage and number of inliers. We propose a novel dense matching module that produces accurate and geometry consistent dense matches.
Given images $I^{1}$ and $I^{2}$, our method can control the memory and computation usage by only selecting the top-K coarse keypoints $p_{c}^{1,2}$ according to their matchability score $M^{1,2}$. Then we can obtain the coarse matches $m_{c}$ using equation Eq. 8, similar to how we obtain the sparse matches $m_{s}$. These matches are naturally coarse matches and require refinement for sub-pixel level accuracy as $M$’s resolution is only $1/4$ of the original resolution. Unlike existing methods [42, 5] that crop fine level features around coarse matches in $I^{1}$ and $I^{2}$ and then predict offset in $I^{2}$, our work proposes a simple, efficient, and accurate module for semi-dense feature matching, inspired by [4] that utilizes sparse correspondences in guiding semi-dense feature matching.
Differing from [4], which iteratively reassigns the coarse matches of detector-free methods, we adopt a similar idea to refine coarse matches to fine matches. With $m_{s}$ obtained from sparse matching, we use the eight-point method to estimate a fundamental matrix $F$ that relates corresponding points between $I^{1}$ and $I^{2}$. We keep $p_{c}^{1}$ unchanged and refine $p_{c}^{2}$ by solving offsets ($\Delta x,\Delta y$) that satisfy the epipolar constraint defined by $F$. First, we convert $p_{c}^{1}$ to homogeneous coordinates $p_{h}^{1}$. Then, epipolar lines $L$ in $I^{2}$ for $p_{h}^{1}$ is computed by
@@ 3.4 Implementation Details
We train RDD with pseudo ground truth correspondences from the MegaDepth dataset [22] and our Air-to-Ground dataset. We collect data from Internet drone videos and ground images of different famous landmarks, more details can be found in the supplementary. Following [42, 38], we use poses and depths to establish pixel correspondences. For every given image pair $(I^{1},I^{2})$ with $N$ matching pixels, $m_{gt}\in R^{N\times 4}$ is defined such that the first two columns are the pixel coordinates of points in $I^{1}$ and the last two columns are those in $I^{2}$ obtained through a warping function $f_{warp}$. We train the descriptor branch and the keypoint branch separately.
@@
To supervise the local feature descriptors, we apply focal loss [23] to focus on challenging examples and reduce the influence of easily classified matches. Given descriptor maps $D^{1}$ and $D^{2}$, we sample descriptor sets $d_{gt}^{1}$ and $d_{gt}^{2}$ based on the ground truth matchability $m_{gt}/k$, where $k$ is the patch size. The $i$-th rows, $d_{gt}^{1}(i)$ and $d_{gt}^{2}(i)$, represent descriptors corresponding to the same points in the images $I^{1}$ and $I^{2}$, respectively. Using these descriptors, we compute a probability matrix $P$ as described in Eq. 8. Our supervision targets positive correspondences only, represented by the diagonal elements $P_{(i,i)}$ of $P$. By enforcing $P_{(i,i)}$ to be close to 1, we calculate the focal loss $\mathcal{L}_{focal}$ as follows:
@@
Figure 3: Qualitative Results on MegaDepth. RDD* and RDD are qualitatively compared to DeDoDe-V2-G [10, 11] ALIKED [50] and XFeat* [33]. RDD* and RDD are more robust compared DeDoDe-G* and ALIKED under challenging scenarios like large scale and viewpoint changes. The red color indicates epipolar error beyond $1\times 10^{-4}$ (in the normalized image coordinates)
@@
The final loss for the descriptor branch is defined as $\mathcal{L}_{D}=\mathcal{L}_{focal}+\mathcal{L}_{matchability}$. We train the descriptor branch on the mixture of MegaDepth dataset [22] and collected Air-to-Ground dataset, with images resized to 800. The training uses a batch size of 32 image pairs and the AdamW optimizer [27]. We use a base learning rate of $1\cdot 10^{-4}$ with a step learning rate scheduler that updates the learning rate every 2000 steps after 20000 steps with $\gamma=0.5$. The descriptor branch converges after 1 day of training on 8 NVIDIA H100 GPU.
@@
We freeze the weights of descriptor branch once it converges and only supervise the keypoint branch to detect accurate and repeatable keypoints. We apply reprojection loss $\mathcal{L}_{reprojection}$, reliability loss $\mathcal{L}_{reliability}$ similar to [49, 8, 28] and dispersity peaky loss $\mathcal{L}_{peaky}$ [49] to train repeatable and reliable keypoints.
@@
Eq. 15 optimizes the scores of keypoints in local window through the soft term $[\hat{i},\hat{j}]_{soft}^{T}$ of Eq. 3, which might affect the first term $[x^{\prime},y^{\prime}]_{nms}^{T}$ of the equation. In order to align their optimization direction, we define the dispersity peak loss similiar to [49] as
@@
$N$ is the window size and $s^{\prime}(i,j)$ are softmax scores defined in Eq. 1.
The overall loss for the keypoint branch is defined as $\mathcal{L}_{K}=\mathcal{L}_{reprojection}+\mathcal{L}_{reliability}+\mathcal{L}_{peaky}$. We use the DKD [49] with a window size of $N=5$ to detect the top 500 keypoints, and we randomly sample 500 keypoints from non-salient positions. We train the keypoint branch similarly to the descriptor branch, the keypoint branch converges after 4 hours of training on an NVIDIA H100 GPU.
@@ 4 Experiments
In this section, we demonstrate the performance of our method in Relative Pose Estimation on public benchmark datasets MegaDepth [22] and collected Air-to-Ground dataset, and in Visual Localization using Aachen Day-Night dataset [39]. All experiments are conducted on an NVIDIA H100 GPU.
@@ 4.1 Relative Pose Estimation
Following [10, 33, 8], we use the standard MegaDepth-1500 benchmark from D2-Net [8] for outdoor pose estimation and adopt the same test split. To further validate the performance of RDD in challenging scenarios, we construct two benchmark datasets. MegaDepth-View is derived from MegaDepth [22] test scenes, focusing on pairs with large viewpoint and scale changes; additionally, we sampled 1500 pairs from our collected Air-to-Ground dataset. This benchmark is designed to validate feature matching performance for large-baseline outdoor cross-view images, which is naturally challenging because of geometric distortion and significant viewpoint changes.
Table 1: SotA comparison on the MegaDepth [22]. Results are measured in AUC (higher is better). Top 4,096 features used to all sparse matching methods. Best in bold, second best underlined.
@@
DKM [9] CVPR’23 60.4 74.9 85.1 67.4 80.0 88.2
RoMa [12] CVPR’24 62.6 76.7 86.3 69.9 81.8 89.4
@@
LoFTR [42] CVPR’21 52.8 69.2 81.2 50.6 65.8 77.9
ASpanFormer [5] ECCV’22 55.3 71.5 83.1 61.3 75.2 84.6
ELoFTR [47] CVPR’24 56.4 72.2 83.5 60.2 74.8 84.7
XFeat* [33] CVPR’24 38.6 56.1 70.5 23.9 37.7 51.8
@@
SP [6]+SG [38] CVPR’19 49.7 67.1 80.6 51.5 65.5 76.9
SP [6]+LG [24] ICCV’23 49.9 67.0 80.1 52.4 67.3 78.5
Dedode-V2-G [10, 11]+LG [24] ICCV’23 44.1 62.1 76.5 41.9 57.1 70.0
RDD+LG [24] ICCV’23 52.3 68.9 81.8 54.2 69.3 80.3
@@
SuperPoint [6] CVPRW’18 24.1 40.0 54.7 7.50 13.3 21.1
DISK [44] NeurIps’20 38.5 53.7 66.6 30.4 41.9 51.6
ALIKED [50] TIM’23 41.8 56.8 69.6 30.0 41.8 53.2
XFeat [33] CVPR’24 24.0 40.1 55.8 8.82 16.5 26.8
DeDoDe-V2-G [10, 11] CVPRW’24, 3DV’24 47.2 63.9 77.5 33.1 47.6 60.2
@@
Method @$5^{\circ}$$10^{\circ}$$20^{\circ}$
SuperPoint [6] CVPRW’18 1.89 3.56 6.81
DISK [44] NeurIps’20 19.2 27.1 34.6
ALIKED [50] TIM’23 12.0 17.8 25.8
XFeat [33] CVPR’24 6.12 11.4 17.5
DeDoDe-V2-G [10, 11] CVPRW’24,3DV’24 31.5 45.3 58.3
@@
We report the AUC of recovered pose under threshold of (5°, 10° ,and 20°). We use RANSAC to estimate the essential matrix. We compared RDD against the state-of-the-art detector/descriptor methods [6, 44, 49, 50, 13, 10, 33] and learned matching methods [38, 24, 42, 5, 47, 9, 12] on MegaDepth [22] dataset. For Air-to-Ground dataset, RDD is compared against detector/descriptor methods. More results on Air-to-Ground data can be found in the supplementary. For all sparse methods, we use resolution whose larger dimension is set to 1,600 pixels and use top 4,096 features, while all other learned matching methods follow experiment settings mentioned in their paper.
@@ Results
The results are presented in Tab. 1 and Tab. 6, and visually in Fig. 3. RDD demonstrates superior performance in the sparse matching setting compared to the state-of-the-art method DeDoDe-V2-G [10, 11], especially in more challenging scenarios. RDD with learned matcher LightGlue [24] also outperforms previous methods.
@@ 4.2 Homography Estimation
We evaluate the quality of correspondences estimated by RDD on HPatches dataset [1] for Homography Estimation. HPatches contains 52 sequences under significant illumination changes and 56 sequences that exhibit large variation in viewpoints. We use RANSAC to robustly estimate the homography given the estimated correspondences. All images are resized such that their shorter size is equal to 480 pixels.
@@
SuperPoint [6] CVPRW’18 93.0 98.0 99.9 70.0 82.0 87.0
DISK [44] NeurIps’20 96.0 98.0 98.0 72.0 81.0 84.0
ALIKED [50] TIM’23 97.0 99.0 99.0 78.0 85.0 88.0
DeDoDe-G [10] 3DV’24 96.0 99.0 99.9 68.0 77.0 80.0
XFeat [33] CVPR’24 90.0 96.0 98.0 55.0 72.0 80.0
@@
We follow [33, 49] to estimate the mean homogeneity accuracy (MHA) with a predefined threshold of {3, 5, 7} pixels. Accuracy is computed using the average corner error in pixels by warping reference image corners onto target images using both ground truth and estimated homographies.
@@ 4.3 Visual Localization
We further validate the performance of RDD on the task of visual localization, which estimates the poses of a given query image with respect to the corresponding 3D scene reconstruction. We evaluate RDD on the Aachen Day-Night dataset [39]. It focuses on localizing high-quality night-time images against a day-time 3D model. There are 14,607 images with changing conditions of weather, season, viewpoints, and day-night cycles.
@@
SuperPoint [6] CVPRW’18 87.4 / 93.2 / 97.0 77.6 / 85.7 / 95.9
DISK [44] NeurIps’20 86.9 / 95.1 / 97.8 83.7 / 89.8 / 99.0
ALIKE [49] TM’22 85.7 / 92.4 / 96.7 81.6 / 88.8 / 99.0
ALIKED [50] TIM’23 86.5 / 93.4 / 96.8 85.7 / 91.8 / 96.9
XFeat [33] CVPR’24 84.7 / 91.5 / 96.5 77.6 / 89.8 / 98.0
@@
[…]

### v5

@@ 3.3 Feature Matching
Dual-Softmax operator [rocco2018neighbourhoodconsensusnetworks, tyszkiewicz2020disk] is used to establish correspondences between two descriptors $d^{1,2}$. The score matrix $S$ between the descriptors is first calculated by $S\left(i,j\right)=\frac{1}{\tau}\cdot\langle\ d^{1},d^{2}\rangle$, where $\tau$ is the temperature. We can now apply softmax on both dimensions of $S$ to obtain the probability of soft mutual nearest neighbor (MNN) matching. The matching probability matrix $P$ is obtained by:
@@
Recent works [sun2021loftr, chen2022aspanformer, Huang2023adamatcher, wang2022matchformer] demonstrated the benefits of semi-dense feature matching by improving coverage and number of inliers. We propose a novel dense matching module that produces accurate and geometry consistent dense matches.
Given images $I^{1}$ and $I^{2}$, our method can control the memory and computation usage by only selecting the top-K coarse keypoints $p_{c}^{1,2}$ according to their matchability score $M^{1,2}$. Then we can obtain the coarse matches $m_{c}$ using equation Eq. 8, similar to how we obtain the sparse matches $m_{s}$. These matches are naturally coarse matches and require refinement for sub-pixel level accuracy as $M$’s resolution is only $1/4$ of the original resolution. Unlike existing methods [sun2021loftr, chen2022aspanformer] that crop fine level features around coarse matches in $I^{1}$ and $I^{2}$ and then predict offset in $I^{2}$, our work proposes a simple, efficient, and accurate module for semi-dense feature matching, inspired by [chen2024geometryawarefeaturematchinglargescale] that utilizes sparse correspondences in guiding semi-dense feature matching.
Differing from [chen2024geometryawarefeaturematchinglargescale], which iteratively reassigns the coarse matches of detector-free methods, we adopt a similar idea to refine coarse matches to fine matches. With $m_{s}$ obtained from sparse matching, we use the eight-point method to estimate a fundamental matrix $F$ that relates corresponding points between $I^{1}$ and $I^{2}$. We keep $p_{c}^{1}$ unchanged and refine $p_{c}^{2}$ by solving offsets ($\Delta x,\Delta y$) that satisfy the epipolar constraint defined by $F$. First, we convert $p_{c}^{1}$ to homogeneous coordinates $p_{h}^{1}$. Then, epipolar lines $L$ in $I^{2}$ for $p_{h}^{1}$ is computed by
@@ 3.4 Implementation Details
We train RDD with pseudo ground truth correspondences from the MegaDepth dataset [li2018megadepth] and our Air-to-Ground dataset. We collect data from Internet drone videos and ground images of different famous landmarks, more details can be found in the supplementary. Following [sun2021loftr, sarlin20superglue], we use poses and depths to establish pixel correspondences. For every given image pair $(I^{1},I^{2})$ with $N$ matching pixels, $m_{gt}\in R^{N\times 4}$ is defined such that the first two columns are the pixel coordinates of points in $I^{1}$ and the last two columns are those in $I^{2}$ obtained through a warping function $f_{warp}$. We train the descriptor branch and the keypoint branch separately.
@@
To supervise the local feature descriptors, we apply focal loss [lin2018focallossdenseobject] to focus on challenging examples and reduce the influence of easily classified matches. Given descriptor maps $D^{1}$ and $D^{2}$, we sample descriptor sets $d_{gt}^{1}$ and $d_{gt}^{2}$ based on the ground truth matchability $m_{gt}/k$, where $k$ is the patch size. The $i$-th rows, $d_{gt}^{1}(i)$ and $d_{gt}^{2}(i)$, represent descriptors corresponding to the same points in the images $I^{1}$ and $I^{2}$, respectively. Using these descriptors, we compute a probability matrix $P$ as described in Eq. 8. Our supervision targets positive correspondences only, represented by the diagonal elements $P_{(i,i)}$ of $P$. By enforcing $P_{(i,i)}$ to be close to 1, we calculate the focal loss $\mathcal{L}_{focal}$ as follows:
@@
Figure 3: Qualitative Results on MegaDepth. RDD* and RDD are qualitatively compared to DeDoDe-V2-G [edstedt2024dedode, edstedt2024dedodev2] ALIKED [zhao2023aliked] and XFeat* [potje2024cvpr]. RDD* and RDD are more robust compared DeDoDe-G* and ALIKED under challenging scenarios like large scale and viewpoint changes. The red color indicates epipolar error beyond $1\times 10^{-4}$ (in the normalized image coordinates)
@@
The final loss for the descriptor branch is defined as $\mathcal{L}_{D}=\mathcal{L}_{focal}+\mathcal{L}_{matchability}$. We train the descriptor branch on the mixture of MegaDepth dataset [li2018megadepth] and collected Air-to-Ground dataset, with images resized to 800. The training uses a batch size of 32 image pairs and the AdamW optimizer [loshchilov2019decoupledweightdecayregularization]. We use a base learning rate of $1\cdot 10^{-4}$ with a step learning rate scheduler that updates the learning rate every 2000 steps after 20000 steps with $\gamma=0.5$. The descriptor branch converges after 1 day of training on 8 NVIDIA H100 GPU.
@@
We freeze the weights of descriptor branch once it converges and only supervise the keypoint branch to detect accurate and repeatable keypoints. We apply reprojection loss $\mathcal{L}_{reprojection}$, reliability loss $\mathcal{L}_{reliability}$ similar to [Zhao2022ALIKE, dusmanu2019d2net, luo2020aslfeat] and dispersity peaky loss $\mathcal{L}_{peaky}$ [Zhao2022ALIKE] to train repeatable and reliable keypoints.
@@
Eq. 15 optimizes the scores of keypoints in local window through the soft term $\hat{i},\hat{j}]_{soft}^{T}$ of Eq. 3, which might affect the first term $x^{\prime},y^{\prime}]_{nms}^{T}$ of the equation. In order to align their optimization direction, we define the dispersity peak loss similiar to [Zhao2022ALIKE] as
@@
NN is the window size and $s^{\prime}(i,j)$ are softmax scores defined in Eq. 1.
The overall loss for the keypoint branch is defined as $\mathcal{L}_{K}=\mathcal{L}_{reprojection}+\mathcal{L}_{reliability}+\mathcal{L}_{peaky}$. We use the DKD [Zhao2022ALIKE] with a window size of $N=5$ to detect the top 500 keypoints, and we randomly sample 500 keypoints from non-salient positions. We train the keypoint branch similarly to the descriptor branch, the keypoint branch converges after 4 hours of training on an NVIDIA H100 GPU.
@@ 4 Experiments
In this section, we demonstrate the performance of our method in Relative Pose Estimation on public benchmark datasets MegaDepth [li2018megadepth] and collected Air-to-Ground dataset, and in Visual Localization using Aachen Day-Night dataset [sattler2018benchmarking6dofoutdoorvisual]. All experiments are conducted on an NVIDIA H100 GPU.
@@ 4.1 Relative Pose Estimation
Following [edstedt2024dedode, potje2024cvpr, dusmanu2019d2net], we use the standard MegaDepth-1500 benchmark from D2-Net [dusmanu2019d2net] for outdoor pose estimation and adopt the same test split. To further validate the performance of RDD in challenging scenarios, we construct two benchmark datasets. MegaDepth-View is derived from MegaDepth [li2018megadepth] test scenes, focusing on pairs with large viewpoint and scale changes; additionally, we sampled 1500 pairs from our collected Air-to-Ground dataset. This benchmark is designed to validate feature matching performance for large-baseline outdoor cross-view images, which is naturally challenging because of geometric distortion and significant viewpoint changes.
Table 1: SotA comparison on the MegaDepth [li2018megadepth]. Results are measured in AUC (higher is better). Top 4,096 features used to all sparse matching methods. Best in bold, second best underlined.
@@
DKM [edstedt2023dkm] CVPR’23 60.4 74.9 85.1 67.4 80.0 88.2
RoMa [edstedt2024roma] CVPR’24 62.6 76.7 86.3 69.9 81.8 89.4
@@
LoFTR [sun2021loftr] CVPR’21 52.8 69.2 81.2 50.6 65.8 77.9
ASpanFormer [chen2022aspanformer] ECCV’22 55.3 71.5 83.1 61.3 75.2 84.6
ELoFTR [wang2024eloftr] CVPR’24 56.4 72.2 83.5 60.2 74.8 84.7
XFeat* [potje2024cvpr] CVPR’24 38.6 56.1 70.5 23.9 37.7 51.8
@@
SP [superpoint2018]+SG [sarlin20superglue] CVPR’19 49.7 67.1 80.6 51.5 65.5 76.9
SP [superpoint2018]+LG [lindenberger2023lightglue] ICCV’23 49.9 67.0 80.1 52.4 67.3 78.5
Dedode-V2-G [edstedt2024dedode, edstedt2024dedodev2]+LG [lindenberger2023lightglue] ICCV’23 44.1 62.1 76.5 41.9 57.1 70.0
RDD+LG [lindenberger2023lightglue] ICCV’23 52.3 68.9 81.8 54.2 69.3 80.3
@@
SuperPoint [superpoint2018] CVPRW’18 24.1 40.0 54.7 7.50 13.3 21.1
DISK [tyszkiewicz2020disk] NeurIps’20 38.5 53.7 66.6 30.4 41.9 51.6
ALIKED [zhao2023aliked] TIM’23 41.8 56.8 69.6 30.0 41.8 53.2
XFeat [potje2024cvpr] CVPR’24 24.0 40.1 55.8 8.82 16.5 26.8
DeDoDe-V2-G [edstedt2024dedode, edstedt2024dedodev2] CVPRW’24, 3DV’24 47.2 63.9 77.5 33.1 47.6 60.2
@@
Method $5^{\circ} @$$10^{\circ} @$$20^{\circ}$
SuperPoint [superpoint2018] CVPRW’18 1.89 3.56 6.81
DISK [tyszkiewicz2020disk] NeurIps’20 19.2 27.1 34.6
ALIKED [zhao2023aliked] TIM’23 12.0 17.8 25.8
XFeat [potje2024cvpr] CVPR’24 6.12 11.4 17.5
DeDoDe-V2-G [edstedt2024dedode, edstedt2024dedodev2] CVPRW’24,3DV’24 31.5 45.3 58.3
@@
We report the AUC of recovered pose under threshold of (5°, 10° ,and 20°). We use RANSAC to estimate the essential matrix. We compared RDD against the state-of-the-art detector/descriptor methods [superpoint2018, tyszkiewicz2020disk, Zhao2022ALIKE, zhao2023aliked, gleize2023silksimplelearned, edstedt2024dedode, potje2024cvpr] and learned matching methods [sarlin20superglue, lindenberger2023lightglue, sun2021loftr, chen2022aspanformer, wang2024eloftr, edstedt2023dkm, edstedt2024roma] on MegaDepth [li2018megadepth] dataset. For Air-to-Ground dataset, RDD is compared against detector/descriptor methods. More results on Air-to-Ground data can be found in the supplementary. For all sparse methods, we use resolution whose larger dimension is set to 1,600 pixels and use top 4,096 features, while all other learned matching methods follow experiment settings mentioned in their paper.
@@ Results
The results are presented in Tab. 1 and Tab. 6, and visually in Fig. 3. RDD demonstrates superior performance in the sparse matching setting compared to the state-of-the-art method DeDoDe-V2-G [edstedt2024dedode, edstedt2024dedodev2], especially in more challenging scenarios. RDD with learned matcher LightGlue [lindenberger2023lightglue] also outperforms previous methods.
@@ 4.2 Homography Estimation
We evaluate the quality of correspondences estimated by RDD on HPatches dataset [hpatches_2017_cvpr] for Homography Estimation. HPatches contains 52 sequences under significant illumination changes and 56 sequences that exhibit large variation in viewpoints. We use RANSAC to robustly estimate the homography given the estimated correspondences. All images are resized such that their shorter size is equal to 480 pixels.
@@
SuperPoint [superpoint2018] CVPRW’18 93.0 98.0 99.9 70.0 82.0 87.0
DISK [tyszkiewicz2020disk] NeurIps’20 96.0 98.0 98.0 72.0 81.0 84.0
ALIKED [zhao2023aliked] TIM’23 97.0 99.0 99.0 78.0 85.0 88.0
DeDoDe-G [edstedt2024dedode] 3DV’24 96.0 99.0 99.9 68.0 77.0 80.0
XFeat [potje2024cvpr] CVPR’24 90.0 96.0 98.0 55.0 72.0 80.0
@@
We follow [potje2024cvpr, Zhao2022ALIKE] to estimate the mean homogeneity accuracy (MHA) with a predefined threshold of {3, 5, 7} pixels. Accuracy is computed using the average corner error in pixels by warping reference image corners onto target images using both ground truth and estimated homographies.
@@ 4.3 Visual Localization
We further validate the performance of RDD on the task of visual localization, which estimates the poses of a given query image with respect to the corresponding 3D scene reconstruction. We evaluate RDD on the Aachen Day-Night dataset [sattler2018benchmarking6dofoutdoorvisual]. It focuses on localizing high-quality night-time images against a day-time 3D model. There are 14,607 images with changing conditions of weather, season, viewpoints, and day-night cycles.
@@
SuperPoint [superpoint2018] CVPRW’18 87.4 / 93.2 / 97.0 77.6 / 85.7 / 95.9
DISK [tyszkiewicz2020disk] NeurIps’20 86.9 / 95.1 / 97.8 83.7 / 89.8 / 99.0
ALIKE [Zhao2022ALIKE] TM’22 85.7 / 92.4 / 96.7 81.6 / 88.8 / 99.0
ALIKED [zhao2023aliked] TIM’23 86.5 / 93.4 / 96.8 85.7 / 91.8 / 96.9
XFeat [potje2024cvpr] CVPR’24 84.7 / 91.5 / 96.5 77.6 / 89.8 / 98.0
@@
Following [potje2024cvpr], HLoc [sarlin2019coarse] is used to evaluate all approaches [potje2024cvpr, Zhao2022ALIKE, superpoint2018, sarlin20superglue, tyszkiewicz2020disk, zhao2023aliked]. We resize the images such that their longer side is equal to 1,024 pixels and sample the top 4,096 keypoints for all approaches. We use the evaluation tool provide by [sattler2018benchmarking6dofoutdoorvisual] to calculate the metrics, the AUC of estimated camera poses under threshold of $\{0.25$m,0.5,0.5m,5,5m}\} for translation errors and $\{2\degree,5\degree,10\degree\}$ for rotation errors respectively.